OpenStudy (anonymous):

What does it mean for an equation to be balanced and why must you keep an equation in balance?

OpenStudy (tkhunny):

Mathematics or Stoichiometry? Balanced: What goes in is what comes out. Unbalanced: Matter and anti-matter mix and the universe is destroyed!

OpenStudy (unklerhaukus):

when two expressions are equal to one another , they can be equated to one another in an equation, \[4=3+1\] and equation is a bit like a set of scales, if you add the same amount of something to both sides do a balanced set of scales they will stay balanced, \[4+10=3+1+10\] also if you subtract the same amount from balanced scales they will remain balanced \[4+10-2=3+1+10-2\] this also works with multiplication \[\frac{4+10-2}4=\frac{3+1+10-2}4\] 

OpenStudy (anonymous):

Thank you c: